OpenStudy (anonymous):

Cansomebody help me with this Solve the following systems of inequalities using the graphing method. Shade the solution set. 17. x+y>4 y+3x<6

OpenStudy (anonymous):

[1] x + y = 4 [2] 3x + y = 6 y = -3x + 6 x + (-3x+6) = 4 -2x = -2 2x = 2 x = 1 // By now we know this much : x = 1 y = -3x+6 Use the x value to solve for y y = -3(1)+6 = 3 {x,y} = {1,3}